OpenStudy (anonymous):

what is the slope of a line that passes through te points (9,4) and (11,-2)

OpenStudy (anonymous):

(4+2)/(9-11) = 6/(-2) = -3

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 9}}\quad ,&{\color{blue}{ 4}})\quad &({\color{red}{ 11}}\quad ,&{\color{blue}{ -2}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)